# Don’t Use LLMs to Make Relevance Judgments

### v1

@@ 6 A New Hope
If the evaluation model has access to privileged information, for example by being fine-tuned on manual relevance judgments on the evaluation topics, then those relevance judgments should still be able to measure systems that use the untuned model. This is what makes Büttcher et al’s [3] results work: the model is trained on relevance data, and so the model is more privileged than any system it is measuring. We actually already knew this: the fact that relevance feedback improves retrieval is a basic result in IR. If we have relevance information gleaned from many systems, essentially a fusion result, the outputs will perform better than any individual system and thus we can measure the individual systems.
This still has the problem that the evaluation isn’t future-proof: we might have a new model that outperforms the relevance feedback of the prior generation.
@@
One simple idea that seems promising is to employ a LLM to follow the assessor and look out for mistakes. This can’t be done by simply asking, “Did you mean ‘relevant’ instead?” since people are primed to trust the computer more than they should. But it may be possible to automate a quality-control process using a model.
We might imagine using LLMs to automate subjects in a user study, or assist researchers coding their observations, or related ideas that might serve to scale up user studies. In these cases, the LLM is not generating the ground truth.
As long as we are not generating ground truth we will use to measure systems, we can use the LLMs to support evaluation. Generating a proxy for ground truth still fails: using relevance judgments to train a model of user satisfaction (or some other evaluation target not directly related to the relevance judgments), and using that model to measure the systems, for example.
@@ 7 Conclusions
This is similar to model collapse [6]. When you train the model using its own outputs, the performance of the model decreases. The collapse mechanism is the measurement error that comes from generating the truth data using the evaluated system. In this case, evaluation is a loss function computed based on the generated truth data.
This doesn’t mean that LLMs can’t allow us to do amazing things. As someone who got his start in IR working with LSI [4], which is essentially an optimal linear embedding, I am very excited by the idea of nonlinear embeddings. IR systems that use LLMs to surmount the vocabulary boundary have enormous promise for real users.
@@
8 Acknowledgments
I thank Ellen Voorhees and Rikiya Takehi for their comments on the talk and on early versions of this paper. I also thank the attendees of the SIGIR 2024 LLM4Eval workshop for their insightful questions.

### v2

@@ 7 Limitations
We actually already knew this: the fact that relevance feedback improves retrieval is a basic result in IR. If we have relevance information gleaned from many systems as we do when pooling, the outputs will perform better than any individual system and thus we can measure any individual systems with less information about relevance than the collective pool.
This still has the problem that the evaluation isn’t future-proof: we might have a new model that outperforms the relevance feedback of the prior generation. We haven’t seen this yet when the collections are pooled from older systems only, if the collection is well-judged (Voorhees et al., 2022). Or new models might have TREC triples (topic, document, relevance) as an explicit component in their training data and be able to make use of that in a retrieval setting.
@@
One simple idea that seems promising is to employ a LLM to follow the assessor and look out for mistakes. This can’t be done by simply asking, “Did you mean ‘relevant’ instead?” since people are primed to trust the computer more than they should (Logg et al., 2019; Bogert et al., 2021). [^8] Logg et al. (2019) is also interesting as much behavioral literature four to five years prior seemed to find that people distrusted algorithmic recommendations. Perhaps our perspectives are changing with exposure. But see also Sparrow et al. (2011) on search’s effects on memory. But it may be possible to automate a quality-control process using a model.
There are evaluation activities that don’t involve creating an answer key. For example, in a user study, researchers observe user behavior and analyze those observations to draw conclusions about the experiment. LLMs might be useful in supporting the observational process (perhaps by transcribing mouse movements and clicks in a readable way) or the analysis process (much as we use statistical models to determine significance).
At the SIGIR workshop, a questioner asked if an LLM-generated evaluation might still be useful even given its flaws. For example, Thomas et al. (2024) found LLM judgments to be as useful as crowdsourced judgments, but not better than curated judgments from a trained team. In their setup, crowd judgments represented a low rung in a tiered hierarchy of relevance judgments and system measurements. If the judgments are not meant to support a rigorous evaluation but rather as noisy training data, then the LLM judgments may be useful. But if the LLM creating the truth data is part of the search system, or is of an older generation than the search system, the results may under-report performance and not be able to distinguish improvements, as shown above. In all cases it should be kept in mind that the ideal used as a comparison point is not human performance, but model performance.
@@ 8 Conclusions
This is similar to model collapse (Guo et al., 2024). When you train the model using its own outputs, the performance of the model decreases. The collapse mechanism is the measurement error that comes from generating the truth data using the evaluated system. In this case, evaluation is a loss function computed based on the generated truth data.
This doesn’t mean that LLMs can’t allow us to do amazing things. As someone who got his start in IR working with LSI (Deerwester et al., 1990), which is essentially an optimal linear embedding, I am very excited by the idea of nonlinear embeddings. IR systems that use LLMs to surmount the vocabulary boundary have enormous promise for real users.
@@
9 Acknowledgments and Disclosure of Funding
I thank Ellen Voorhees and Rikiya Takehi for their comments on the talk and on early versions of this paper. I also thank the attendees of the SIGIR 2024 LLM4Eval workshop for their insightful questions and continuing discussions.
No external funding was received in support of this work.
The TREC activity has been annually reviewed by NIST’s Research Protection Office and determined to not be human subjects research.
Any company, product or service mentioned in this paper should not be taken as an endorsement of that company, product, or service by NIST. Nothing in this paper should be read as a comparison to or among commercial products.